evaluate each algebraic expression for the given value of the variable. -x³ - x when x = -2
Please show steps!!

Answer by edjones(7569)

-x³ - x
sorry, I didn't notice that x=-2
-(-2*-2*-2)-(-2)=-(-8)+2=+8+2=10
THIS IS MY FINAL ANSWER
Ed